OpenStudy (anonymous):

Find the common ratio of the following geometric sequence. If necessary, use the slash bar ( / ) to enter a fraction. Reduce fractions to their lowest terms. (5/12),(1/3),(4/15),(16/75),(64/375)

OpenStudy (a_clan):

4/5

OpenStudy (anonymous):

r = 1/3 / 5/12 = 1/3 * 12/5 = 12/15 = 4/5